OpenStudy (anonymous):

Help.. ;o –3(5 – 15n) = 15(3n – 2)

OpenStudy (zzr0ck3r):

distribute the \(3\) on the left and the \(15\) on the right and what do you have?

OpenStudy (anonymous):

–15 and –45

OpenStudy (anonymous):

Distribute the -3 and distribute 15.

OpenStudy (anonymous):

;o

OpenStudy (anonymous):

Corect! For the first equation.

OpenStudy (anonymous):

It equals 30? o;

OpenStudy (muzzack):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : \(\cal\large\color{lime}{-3*(5-15*n)-(15*(3*n-2))=0}~\) Step 1 : Multiply \(\huge\cal\color{lime}{3~by~5-15n~}\) Pull out like factors : \(\huge\cal\color{violet}{5 - 15n = -5 • (3n - 1)~}\) Equation at the end of step 1 : \(\large\color{blue}{(0 - -15 • (3n - 1)) - 15 • (3n - 2) = 0~}\) Step 2 : Simplify 15•(3n-1) - 15•(3n-2) Pulling out like terms : Pull out 15 After pulling out, we are left with : \(\large\cal\color{green}{15 • ( 1 * (3n-1) +( (-1) * (3n-2) ))~}\) Equation at the end of step 2 : 15 = 0 Step 3 : Solve 15 = 0 Equations which are never true : Solve : 15 = 0 This equation has no solution. A a non-zero constant never equals zero. No Solutions found

OpenStudy (muzzack):

so there is no solution @yourmellody

OpenStudy (anonymous):

Thank u :33

OpenStudy (muzzack):

\(\huge\cal\color{green}{your~very~welcome}\)